OpenStudy (anonymous):

What is the limit of x - x^2 * ln((1 + x) / x) when x approaches infinity?

OpenStudy (zarkon):

\[\lim_{x\to\infty}\left[x - x^2 \cdot \ln\left(\frac{1 + x}{x}\right)\right]\] Let \(\displaystyle u=\ln\left(\frac{1 + x}{x}\right)=\ln\left(\frac{1}{x}+1\right)\). Then \(\displaystyle x=\frac{1}{e^u-1}\). Our limit becomes \[\lim_{u\to 0^+}\left[\frac{1}{e^u-1}-\frac{1}{(e^u-1)^2}u\right]\] \[=\lim_{u\to 0^+}\frac{e^u-1-u}{(e^u-1)^2}\] Using L'Hospital's rule yields \[=\lim_{u\to 0^+}\frac{e^u-1}{2(e^u-1)e^u}=\lim_{u\to 0^+}\frac{e^u-1}{2e^{2u}-2e^u}\] Using L'Hospital's rule again \[=\lim_{u\to 0^+}\frac{e^u}{4e^{2u}-2e^u}=\frac{1}{4-2}=\frac{1}{2}\]

OpenStudy (anonymous):

Thanks :)